OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (abhisar):

Hi ! Have you tried to solve it on your own first?

OpenStudy (anonymous):

there is no solution

OpenStudy (abhisar):

-3(x + 3) = -3(x + 1) - 5 \(\Rightarrow\) -3x-9 = =3x-3 - 5

OpenStudy (abhisar):

Now can you solve it?

OpenStudy (anonymous):

no it is -3x-9=-3x-3-5

OpenStudy (anonymous):

I tried, but I am no good at math, I got confused.

OpenStudy (anonymous):

there is no solution

OpenStudy (anonymous):

3x-3x=0 nothing there is no solution